CARA Telecon Minutes Date: March 12, 2001 Time: 1:00PM CDT Attendance: Boston University: Cornell: CMU: J. Peterson Northwestern G. Novak R.I.T. S.A.O A. Lane, T. Stark South Pole: C. Martin, C. Kaminski Chicago: J. Carlstom, A.Harper, S. Meyer, J. Bausch, R. Landsberg, R. Loewenstein, N. Odalen, R. Pernic, UCB: UCSB: J. Ruhl Minutes from last two telecons were approved. Agenda Items News from the Pole: Liquid Helium - C. Kaminski reported that A. Baker stated that the Wessington dewar with the pressure reduced is losing approx. 20 liters per day. The mass and helium level sensors do not yet agree over the short term since the last depressurization. Based on how long things have taken to settle and agree in the past after pressure changes, we think it will be next week before we have a better idea of how the latest modification affected things. J. Ruhl feels something is wrong with the level sensor. J. Ruhl stated that the flow is 20 liters per day and the mass loss shows 25 liters per day so that is good news. There may be an improvement, but still to early to tell. In response to question from S. Meyer -- C. Kaminski stated that there is no flow coming out of the 7 PSIG popper. A. Baker proposed doing the same thing on the Wessington #1 and John Ruhl agreed it was a reasonable next step. J. Ruhl is of the opinion that the Wessington's will not come down to 10 liters per day. Raytheon is actively involved in decision regarding the actions being taken on the dewars. If the Wessington's are losing 20 liters per day then CARA will need to figure out what this means for science this winter. The 3000 gallon dewar level sensor show 3/4 full. Current boil off rate is nominal - estimate 20 liters per day according to A. Barker. *** Action item - C. Kaminski will check on boil off rate and report before the Projects Meeting on Thursday 3/15/01. J. Ruhl wants to know absolute pressure and what how much was filled at McMurdo. Giles Novak has requested a full report on the 3000 Gallon from J. Alcorta, from the time it was filled in McMurdo until it was installed back at the Pole. Giles, Tony and John to come up with options on how to use remain helium before the Projects Meeting on Thursday. J. Carlstrom is concerned that the helium discussions, although a very important issue, is taking up all of the telecon time on Monday. Perhaps some of this discussion could be done during special helium telecons. Projects Meeting planning - - G. Novak will give talk on SPARO - R. Pernic will give report on AASTO - No report from the IR group.(in fact, Al Harper presented Abu data) - T. Stark will bring plots on the FTS - R. Loewenstein will give talk on communications. - R. Pernic commented that if everyone would send their information as promised then he will be able to give presentation. Right now R. Pernic feels that CARA we be over extended. R. Pernic will not have any spread sheets, but will present his understanding of the situation. - AST/RO - will have no new instruments next year at Pole. - John Storey wants Carlstrom/Meyer to push NSF for move of the AASTO. - NSF has not made any commitment on continuation proposals. - CARA will be doing SIPs for the group again. We do not want to have individual projects doing their own SIPs for the upcoming Austral summer. Projects Update - DASI - working, things are fine, doing tests, waiting for winter. ACBAR - still observing. AST/RO - observing, CCD camera exhibits strange behavior. Tipper running, receiver going great.